THE DISTANCE BETWEEN THE PLANE x + 2y - 2z = 3 and 2x + 4y - 4z = 7
Answer by nabla(475)

Note first that they are parallel planes (necessary).
When y=x=0 in x+2y-2z=3, there exists the point (0,0,-3/2)
D=|-4*(-3/2)-7|/sqrt(4+32)=1/6

It should be clear I skipped a few steps in showing the work, I assume you are able to do it. For more information see:
http://www.math.ucla.edu/~ronmiech/Calculus_Problems/32A/chap11/section5/718d65/718_65.html